Two cars are travelling at constant speeds. A Neon is travelling at 15.6 m/s and a Corvette is travelling at 23.9 m/s. When a stopwatch is started, the Corvette is 13.5 meters behind the Neon. How long does it take the Corvette to catch the Neon? How far did the Neon travel before it was overtaken?


1
Expert's answer
2021-09-14T10:59:59-0400

Let "t" be the time Corvette took to catch Neon.

Distance"=rate \u00d7time"

Let position be represented by "x"

Distance position of Corvette is:

"x_{c}=23.9t.....i)"

Position of Neon

"x_{n}=13.5+15.6t.....ii)"

Using i) and ii)

"23.9t=13.5+15.6t"

"(23.9-15.6)t=13.5"

"8.3t=13.5"

"\frac{8.3}{8.3}t=\frac{13.5}{8.3}"

"t=1.63s"


Distance traveled by Neon

"D=15.6\u00d71.63"

"=25.428m"